A bag contains coins of denomination 50 paise and 5 ruppes . The total value of these coins is 300 ruppes . If the total number of coins in the bag is 150 , find the number of coins of each type?? Not able to answer this question . Please help !
THANKS

Grade:7

1 Answers

srikar
9 years ago
let x be the no. of 5 Rs coins
let y be the no. of 0.5 Rs coins
Given,
x+y = 150
and 5x + 0.5 y = 300  =====>10x + y = 600
                                                x + y = 150
on subracting we get 9x + 0 = 450 ==>  x= 50 ==> y= 100
Therefore no of 5 Rs coins is 50  and   0.5 Rs coins is 100